A cessna aircraft has a lift-off speed of 120km/h. (a) what minimum constant acceleration does the aircraft require if it is to be airborne after a takeoff run of 240 m? (b) how long does it take the aircraft to become airborne?
1
Expert's answer
2012-09-18T11:24:55-0400
(a)
As V0 = 0, then the equation of the speed is the following:
V =
at,
and the equation of the covered distance:
S = at^2 / 2
where a -
acceleration.

V = 120 km/h = 100 / 3 m/s.
After multiplying the first
equation by t and division by the second equation we get:

V^2 / S =
2a,
a = V^2 / (2S) = 10000 / (9 * 2 * 240) = 2.31 m / s^2.

(b)
From
the first equation:

t = V / a = 100 / (3 * 2.31) = 14.4 s.